# Light Control of Triplet Pairing in Correlated Electrons with Mixed-Sign Interactions

### v2

@@ II.2 Floquet Engineering of Anisotropic Spin Interactions
Beyond the dominant single-photon resonance, Floquet theory also predicts a two-photon resonance condition, occurring when the driving frequency satisfies $\Omega\approx E_{\mathrm{int}}/2=4.5\,t_{h}$. Variations in local occupations further modulate $E_{\mathrm{int}}$, yielding a relatively narrower yet finite resonant window for the two-photon process, spanning from $3.75t_{h}$ to $5.25t_{h}$. The frequency dependence observed in the simulated $\Phi_{x}(t)$ (Fig. 3) also aligns well with this prediction, further corroborating the role of Floquet engineering in modulating pairing correlations. Additional details on the pairing correlation dynamics at $\Omega=5.2t_{h}$ are provided in Supplementary Note III.
@@ II.3 Ultrafast Switching Between Different Channels
The ability to selectively enhance $p$-wave superconductivity along a given direction raises a compelling question: can optical excitation be used to switch the dominant pairing channel to an orthogonal one? Within a mean-field framework, transitions between different pairing channels like $p_{x}$ and $p_{y}$ (or between chiral states $p_{x}+ip_{y}$ and $p_{x}-ip_{y}$) are inherently challenging due to the strict orthogonality of their order parameters. Typically, such a switch requires a carefully designed sequence of linear and circular pulses to navigate a path on the Bloch sphere that connects the two poles [41, 57]. However, in strongly correlated many-body systems, quantum fluctuations naturally broaden the distribution of pairing correlations across the Bloch sphere rather than confining them to discrete symmetry-breaking points. This broader distribution may allow for more flexible control pathways, enabling direct switching between orthogonal $p$-wave channels without requiring finely tuned orchestrated pulse sequences.
@@
Figure 5a presents the evolution of spin-triplet pairing correlations under the influence of two sequential laser pulses. The first laser pulse, $x$-polarized with an amplitude $A_{0}=0.4(\frac{e}{\hbar c})^{3/2}$ and frequency $\Omega=9.2\,t_{h}$ (same as Fig. 2), triggers the symmetry breaking and enhances the $p_{x}$ pairing correlations. After several oscillation cycles, significantly exceeding the pulse duration, a second $y$-polarized laser pulse is applied at a time delay $\Delta t=25\,t_{h}^{-1}$, maintaining the same pump frequency and phase. To ensure an active switching of the dominant pairing channel, rather than merely canceling the pre-established correlations (see Supplementary Note IV for detailed discussions), the second pulse amplitude is set to $A_{0}^{\prime}=2A_{0}$ (see Fig. 5a). Remarkably, this second pulse rapidly reverses the dominance between $\Phi_{x}$ and $\Phi_{y}$. The $p_{x}$ pairing correlation experiences a substantial suppression, with its average intensity reduced by roughly 43%, while the $p_{y}$ pairing correlation remains relatively less affected, due to the interplay between light-enhanced pairing effects and additional photo-induced carriers.
@@ III Discussion
Our simulations are based on the exact solutions of correlated electrons confined to a finite cluster, where spontaneous symmetry breaking does not occur. As a result, our analysis focuses on short-range pairing correlations rather than off-diagonal long-range orders, which would require extrapolation to the 2D thermodynamic limit — a challenge beyond the capabilities of existing many-body computational techniques. Despite this limitation, short-range quantum fluctuations are a ubiquitous feature of strongly correlated systems. At equilibrium, fluctuation-driven non-symmetry-breaking gaps have been experimentally observed in unconventional superconductors, such as cuprates [58, 59] and FeSe [60, 61], as well as in other phase transitions [62, 63, 64]. These fluctuations become even more pronounced in low-dimensional or nanoscale materials. More importantly, light-induced superconductivity exhibits even stronger quantum fluctuations with shorter-range correlations [65, 66, 51]. These short-range correlations manifest as superconducting gaps with limited coherence, making them effectively indistinguishable from long-range orders in pump-probe spectroscopic measurements.
In strongly correlated systems, the inherent flexibility of pairing correlations offers significant advantages for laser-based control compared, beyond the constraints of mean-field order parameters. As demonstrated above, spin-triplet pairing correlations can be effectively engineered through Floquet manipulation of spin-exchange interactions along the pump polarization. Furthermore, the distributed nature of these correlations across the Bloch sphere enables a straightforward single-pulse strategy to switch between orthogonal pairing channels, which would otherwise necessitate more sophisticated control protocols in a mean-field framework [41, 57]. These fundamental differences suggest that many-body quantum fluctuations, rather than acting as a limiting factor, can instead serve as a powerful resource for the dynamic control of spin-triplet superconducting states. Such a distinction may also account for the fact that light-induced superconductivity has been frequently observed in correlated materials [43, 44, 45, 46, 47, 48].
@@ IV.1 Extended Hubbard Model
where $c_{i\sigma}(c^{\dagger}_{i\sigma})$ annihilates (creates) an electron at site $i$ with spin $\sigma=\uparrow,\downarrow$ and $n_{i\sigma}=c^{\dagger}_{i\sigma}c_{i\sigma}(n_{i}=n_{i\uparrow}+n_{i\downarrow})$ represents the electron density operator. The model incorporates three key parameters: the nearest-neighbor hopping amplitude $t_{h}$, the on-site Coulomb repulsion $U$, and the nearest-neighbor interaction $V$. In this study, the on-site $U$ is always repulsive, accounting for the double-occupation energy, while the nearest-neighbor interaction $V$ is attractive, representing an effective phonon-mediated interaction. We adopt model parameters $U=8t_{h}$ and $V=-t_{h}$, based on prior experimental estimates for 1D cuprate chains [32, 36]. For $t_{h}=300-400\,$meV, a 400-nm laser resides in the single-photon resonance and an 800-nm laser resides in the two-photon resonance. Simulations are conducted on a $\times4$ cluster with periodic boundary conditions.
@@ IV.2 Peierls substitution and the Krylov-Subspace Method
To describe the light-driven dynamics, we incorporate the Peierls substitution $t_{h}c^{\dagger}_{i\sigma}c_{j\sigma}\xrightarrow[]{}t_{h}e^{i\int_{r_{i}}^{r_{j}}{\textbf{A}(t)d\textbf{r}}}c^{\dagger}_{i\sigma}c_{j\sigma}$, which serves as an effective approximation of the light-matter interaction within the second-quantization framework. The specific pump vector potential $\textbf{A}(t)$ is provided by Eqs. (3) and (5) of the main text. We employ the SLEPc library to solve the ground state  [67], which yields a three-fold degenerate ground state under the chosen parameters. The time evolution of the ensemble-averaged density matrix is computed using the Krylov subspace method [68].
@@ Acknowledgements
We acknowledge insightful discussions from Matteo Mitrano. This work is supported by the Air Force Office of Scientific Research Young Investigator Program under grant FA9550-23-1-0153. The simulations presented in this paper were performed on the Frontera computation system at Texas Advanced Computing Center, operated under NSF Award No.OAC-1818253.

### v3

@@ II.2 Floquet Engineering of Anisotropic Spin Interactions
Beyond the dominant single-photon resonance, Floquet theory also predicts a two-photon resonance condition, occurring when the driving frequency satisfies $\Omega\approx E_{\mathrm{int}}/2=4.5\,t_{h}$. Variations in local occupations further modulate $E_{\mathrm{int}}$, yielding a relatively narrower yet finite resonant window for the two-photon process, spanning from $3.75t_{h}$ to $5.25t_{h}$. The frequency dependence observed in the simulated $\Phi_{x}(t)$ (Fig. 3) also aligns well with this prediction, further corroborating the role of Floquet engineering in modulating pairing correlations. Additional details on the pairing correlation dynamics at $\Omega=5.2t_{h}$ are provided in Supplementary Note 3.
Beyond the Floquet-engineered spin exchange $J_{x}^{\rm eff}$, the laser field also modifies the electronic hopping amplitude as $t_{x}^{\rm eff}\sim t\mathcal{J}_{0}(A_{0}\cdot\hat{\mathbf{e}}_{\alpha})$, thereby inducing anisotropy between the $x$ and $y$ directions. Although this anisotropy resembles the effect of uniaxial strain and does contribute to the spin-triplet pairing,[gassner2024light], it does not generate the mixed AFM–FM exchange interactions responsible for the resonances observed in Fig.3, and thus does not serve as the primary mechanism behind the light-induced superconductivity in the EHM. Nevertheless, the suppression of $t_{x}^{\rm eff}$ constrains the strength of light-enhanced pairing, which explains the observed saturation and eventual decline of $\Phi_{x}$ around $A_{0}\sim 1$ in Figs. 3a and b.
@@ II.3 Ultrafast Switching Between Different Channels
The ability to selectively enhance $p$-wave superconductivity along a given direction raises a compelling question: can optical excitation be used to switch the dominant pairing channel to an orthogonal one? Within a mean-field framework, transitions between different pairing channels like $p_{x}$ and $p_{y}$ (or between chiral states $p_{x}+ip_{y}$ and $p_{x}-ip_{y}$) are inherently challenging due to the strict orthogonality of their order parameters. Typically, such a switch requires a carefully designed sequence of linear and circular pulses to navigate a path on the Bloch sphere that connects the two poles [claassen2019universal, yu2021optical]. However, in strongly correlated many-body systems, quantum fluctuations naturally broaden the distribution of pairing correlations across the Bloch sphere rather than confining them to discrete symmetry-breaking points. This broader distribution may allow for more flexible control pathways, enabling direct switching between orthogonal $p$-wave channels without requiring finely tuned orchestrated pulse sequences.
@@
Figure 5a presents the evolution of spin-triplet pairing correlations under the influence of two sequential laser pulses. The first laser pulse, $x$-polarized with an amplitude $A_{0}=0.4$ and frequency $\Omega=9.2\,t_{h}$ (same as Fig. 2), triggers the symmetry breaking and enhances the $p_{x}$ pairing correlations. After several oscillation cycles, significantly exceeding the pulse duration, a second $y$-polarized laser pulse is applied at a time delay $\Delta t=25\,t_{h}^{-1}$, maintaining the same pump frequency and phase. To ensure an active switching of the dominant pairing channel, rather than merely canceling the pre-established correlations (see Supplementary Note 4 for detailed discussions), the second pulse amplitude is set to $A_{0}^{\prime}=2A_{0}$ (see Fig. 5a). Remarkably, this second pulse rapidly reverses the dominance between $\Phi_{x}$ and $\Phi_{y}$. The $p_{x}$ pairing correlation experiences a substantial suppression, with its average intensity reduced by roughly 43%, while the $p_{y}$ pairing correlation remains relatively less affected, due to the interplay between light-enhanced pairing effects and additional photo-induced carriers.
@@ III Discussion
Because of the need to preserve $C_{4}$ symmetry and the inherent numerical difficulties of simulating nonequilibrium dynamics, our analysis is based on exact diagonalization of correlated electrons within a finite cluster, where spontaneous symmetry breaking cannot occur. Accordingly, we focus on short-range pairing correlations rather than off-diagonal long-range order, whose identification would require extrapolation to the 2D thermodynamic limit — a task currently beyond the reach of existing many-body computational techniques. Despite this limitation, short-range quantum fluctuations are a ubiquitous feature of strongly correlated systems. At equilibrium, fluctuation-driven non-symmetry-breaking gaps have been experimentally observed in unconventional superconductors, such as cuprates [kondo2015point, he2021superconducting] and FeSe [faeth2021incoherent, xu2021spectroscopic], as well as in other phase transitions [schafer2001high, chatterjee2015emergence, chen2023role]. These fluctuations become even more pronounced in low-dimensional or nanoscale materials. More importantly, light-induced superconductivity exhibits even stronger quantum fluctuations with shorter-range correlations [boschini2018collapse, sun2020transient, wang2021fluctuating]. These short-range correlations manifest as superconducting gaps with limited coherence, making them effectively indistinguishable from long-range orders in pump-probe spectroscopic measurements.
In strongly correlated systems, the inherent flexibility of pairing correlations offers significant advantages for laser-based control compared, beyond the constraints of mean-field order parameters. As demonstrated above, spin-triplet pairing correlations can be effectively engineered through Floquet manipulation of spin-exchange interactions $J$ along the pump polarization. Furthermore, the distributed nature of these correlations across the Bloch sphere enables a straightforward single-pulse strategy to switch between orthogonal pairing channels, which would otherwise necessitate more sophisticated control protocols in a mean-field framework [claassen2019universal, yu2021optical], or rely on static uniaxial strain or Jahn–Teller distortions [gassner2024light]. These fundamental differences suggest that many-body quantum fluctuations, rather than acting as a limiting factor, can instead serve as a powerful resource for the dynamic control of spin-triplet superconducting states. Such a distinction may also account for the fact that light-induced superconductivity has been frequently observed in correlated materials [fausti2011light, nicoletti2014optically, hu2014optically, mankowsky2014nonlinear, buzzi2020photomolecular, buzzi2021phase].
Our study focuses on the ultrafast regime soon after the pump finishes. Based on estimated hopping parameters typical of cuprates (see Methods), the time interval shown in Figs. 2 and 5 spans approximately $\sim 100$ fs. This short interval lies within the prethermal regime, as identified in related systems [aoki2014nonequilibrium, de2021colloquium]. If the simulation time were extended substantially, it would be necessary to account for additional effects such as heating, decoherence, and the retarded response of phonons.
@@ IV.1 Extended Hubbard Model
where $c_{i\sigma}(c^{\dagger}_{i\sigma})$ annihilates (creates) an electron at site $i$ with spin $\sigma=\uparrow,\downarrow$ and $n_{i\sigma}=c^{\dagger}_{i\sigma}c_{i\sigma}(n_{i}=n_{i\uparrow}+n_{i\downarrow})$ represents the electron density operator. The model incorporates three key parameters: the nearest-neighbor hopping amplitude $t_{h}$, the on-site Coulomb repulsion $U$, and the nearest-neighbor interaction $V$. In this study, the on-site $U$ is always repulsive, accounting for the double-occupation energy, while the nearest-neighbor interaction $V$ is attractive, representing an effective phonon-mediated interaction. We adopt model parameters $U=8t_{h}$ and $V=-t_{h}$, based on prior experimental estimates for 1D cuprate chains [chen2021anomalously, wang2021phonon], while the conclusion broadly applies to a variety of model parameters (see Supplementary Note 5). The simulation is performed within a canonical ensemble framework, where the total particle number is fixed to ensure quarter filling, and the temperature is set to absolute zero. For $t_{h}=300-400\,$meV, a 400-nm laser resides in the single-photon resonance and an 800-nm laser resides in the two-photon resonance. Simulations are conducted on a $\times4$ cluster with periodic boundary conditions.
@@ IV.2 Peierls substitution and the Krylov-Subspace Method
To describe the light-driven dynamics, we incorporate the Peierls substitution $t_{h}c^{\dagger}_{i\sigma}c_{j\sigma}\xrightarrow[]{}t_{h}e^{i\int_{r_{i}}^{r_{j}}{\textbf{A}(t)d\textbf{r}}}c^{\dagger}_{i\sigma}c_{j\sigma}$, which serves as an effective approximation of the light-matter interaction within the second-quantization framework. The specific pump vector potential $\textbf{A}(t)$ is provided by Eqs. (3) and (5) of the main text. We employ the SLEPc library to solve the ground state  [hernandez2005slepc], which yields a three-fold degenerate ground state due to the hidden hypercubic symmetry of the $4\times 4$ cluster. The expectation value of pairing correlations $\Phi_{\alpha}$ is taken as an average over all degenerate states and all triplet $S_{z}$ states at the instantaneous wavefunction $|\psi(t)\rangle$. The time evolution of the ensemble-averaged density matrix is computed using the Krylov subspace method [manmana2007strongly].
@@ Acknowledgements
We acknowledge insightful discussions from Cheng-Chien Chen, Martin Claassen, and Matteo Mitrano. This work is supported by the Air Force Office of Scientific Research Young Investigator Program under grant FA9550-23-1-0153. The simulations presented in this paper were performed on the Frontera computation system at Texas Advanced Computing Center, operated under NSF Award No.OAC-1818253.